OpenStudy (anonymous):

1/4(x-12)+3/4=x/3 Solve for X

OpenStudy (anonymous):

1/4x-3+3/4=x/3step1

OpenStudy (anonymous):

step 2? is?

OpenStudy (anonymous):

x=-27

OpenStudy (anonymous):

-3+3/4=x/3-x/4step2

OpenStudy (anonymous):

im sure you can proccede from there until you get x=-27

OpenStudy (anonymous):

yes thank you.